Grade 11Physical Chemistry

14.7 g of sulphuric acid was needed to dissolve 16.8 g of of metal. Calculate the equivalent weight of the metal and the volume of H2 gas liberated at STP?
Please explain the solution of the above question clearly with steps.
THANKS.

AB
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
moles of H2 would be that of moles of H2 in H2SO4.
Therefore,
moles = 14.7/98 = 0.15
V = 0.15*22.4 = > 3.36 litre.
 
For the eq. weight of metal,
Eq. Weight = moles x ve;cne factor => 0.15 * 2 => 0.3 g